OpenStudy (anonymous):

Indicate which property is illustrated below. A. distributive B. commutative C. identity D. associative

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Step 1 12x + 4 + 5x + 5 = (12x + 4) + (5x + 5) Step 2 = 12x + (4 + 5x) + 5 Step 3 = 12x + (5x + 4) + 5 Step 4 = (12x + 5x) + (4 + 5) Step 5 = (12 + 5)x + (4 + 5) Step 6 = 17x + 9

OpenStudy (anonymous):

which step do you mean?

OpenStudy (anonymous):

sorry, I have to go.

OpenStudy (anonymous):

Indicate which property is illustrated below. A. distributive B. commutative C. identity D. associative

OpenStudy (anonymous):

3x + 8y = 8y + 3x